OpenStudy (anonymous):

Solve the exponential equation. Express your solutions in exact form only. Please show all of your work. 7e^(xlog4) = 2

OpenStudy (mertsj):

\[e ^{xlog4}=\frac{2}{7}\]

OpenStudy (mertsj):

Take the natural log of both sides.

OpenStudy (mertsj):

\[xlog4(\ln e)=\ln \frac{2}{7}\]

OpenStudy (mertsj):

\[x(.602)=-1.253\]